A link is shared with 10 computers, each sending 1 packet/s. The link can serve 10 packets/s. What is the delay?
Answer #1

Number of packets that can be served by the link per second = 10.

Number of connected computers = 10

Number of packet sent from each computer per second = 1

Total number of sent packets in the network per second = (Number of connected computers * Number of packet sent from each computer per second) = 10 * 1 = 10

Utilization of the link = ((10/10) * 100)% = 100%

So, there will be no delay.
Hope this helps.